OpenStudy (banjoman):

MEDAL GIVEN !A rectangular room is 1.9 times as long as it is wide, and its perimeter is 29 meters. Find the dimension of the room. Ive gotten several methods from others on this site, but none have helped me to get the correct answer.

OpenStudy (princeharryyy):

2*(w+1.9w ) =29 2*2.9w =29 2w = 10 w=5 l =1.9*w = 1.9*5 =9.5

OpenStudy (princeharryyy):

2*(l+b) = perimeter of rectangle

OpenStudy (banjoman):

so it would be 2*(9.5+5)?

OpenStudy (princeharryyy):

no. length is 9.5 and width is 5. that's what they are asking for.

OpenStudy (banjoman):

thank you much I'm going to practice more on problems like this, thank you for your help

OpenStudy (princeharryyy):

:)